OpenStudy (anonymous):

eval the fxn at the midpoints f(x)=6/x on [1/2, 11/2] divided into 5 subintervals

OpenStudy (anonymous):

your interval has length \(5\) so dividing it into five sum intervals is straight forward, make each interval length 1, so they are \([\frac{1}{2},\frac{3}{2}],[\frac{3}{2},\frac{5}{2}],...\)

OpenStudy (anonymous):

the midpoint of each interval conveniently is at the integers 1, 2, 3, 4, 5

OpenStudy (anonymous):

then your final job is to compute \\[f(1)+f(2)+f(3)+f(4)+f(5)\]

OpenStudy (anonymous):

the answer is 13.7 but how come the solution is 137/50